OpenStudy (anonymous):

Someone Please Help! Solve a system of linear equations using substitution, elimination OR substitution! 3x - y = -137 y = 2x + 99

OpenStudy (anonymous):

3x - y = -137 y = 2x + 99 3x - y = -137 plug in y from the 2nd equation 3x - (2x + 99) = -137 3x - 2x - 99 = -137 x - 99 = -137 x = -137 + 99 x = -38 y = 2x + 99 y = 2(-38) + 99 y = -76 + 99 y = 23

OpenStudy (anonymous):

Theres also, 2.) -x + 2y = -2 x + 4y = -10 3.) 2x + y = -1 3x - y = 16 I have to solve one for each system of linear equations!:(

OpenStudy (anonymous):

@surdawi did you use substitution, elimination or graphing??

OpenStudy (anonymous):

substitution

OpenStudy (anonymous):

@jazy can show you the other 2 :D

OpenStudy (anonymous):

@surdawi Thank you for the help!! i appreciate it!!:)